40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sun Sep 16 16:42:16 2012, tara, DailyProgress, fiber optic, setup for fiber optic distribution. 2012_09_16_fiber_setup.jpg
    Reply  Mon Sep 17 20:59:15 2012, tara, DailyProgress, fiber optic, setup for fiber optic distribution. 2012_09_18_fiber.jpgIMG_1761.jpg
Message ID: 1047     Entry time: Sun Sep 16 16:42:16 2012     Reply to this: 1049
Author: tara 
Type: DailyProgress 
Category: fiber optic 
Subject: setup for fiber optic distribution. 

I'm setting up fiber optic so that I can send frequency stabilized laser to ATF. Right now the power coming out is small (0.8 mW from 20mW input). I'm working on better mode match for better efficiency.

2012_09_16_fiber_setup.jpg

 

Note: due to the space limitation, I cannot pick the beam after the broadband EOM used for frequency stabilized the laser. The beam is taken after the PMC, where it was used to dumped the excess power.

Collimator: http://www.thorlabs.com/thorProduct.cfm?partNumber=CFC-2X-C

optical fiber: nufern pm980

modematching: The focal length of the collimator is 2.0 mm, MFD of the fiber is 8 um. The beam diameter at the lens is then ~360 um.

The coupled power is quite small. I'll check the mode matching again to get more power coming out.

 

ELOG V3.1.3-